# Rep3D: Re-parameterize Large 3D Kernels with Low-Rank Receptive Modeling for Medical Imaging

### v1

@@ Abstract
In contrast to vision transformers, which model long-range dependencies through global self-attention, large kernel convolutions provide a more efficient and scalable alternative, particularly in high-resolution 3D volumetric settings. However, naïvely increasing kernel size often leads to optimization instability and degradation in performance. Motivated by the spatial bias observed in effective receptive fields (ERFs), we hypothesize that different kernel elements converge at variable rates during training. To support this, we derive a theoretical connection between element-wise gradients and first-order optimization, showing that structurally re-parameterized convolution blocks inherently induce spatially varying learning rates. Building on this insight, we introduce Rep3D, a 3D convolutional framework that incorporates a learnable spatial prior into large kernel training. A lightweight two-stage modulation network generates a receptive-biased scaling mask, adaptively re-weighting kernel updates and enabling local-to-global convergence behavior. Rep3D adopts a plain encoder design with large depthwise convolutions, avoiding the architectural complexity of multi-branch compositions. We evaluate Rep3D on five challenging 3D segmentation benchmarks and demonstrate consistent improvements over state-of-the-art baselines, including transformer-based and fixed-prior re-parameterization methods. By unifying spatial inductive bias with optimization-aware learning, Rep3D offers an interpretable, and scalable solution for 3D medical image analysis. The source code is publicly available at https://github.com/leeh43/Rep3D.
@@ 1 Introduction
The landscape of medical vision models has evolved rapidly, expanding from early convolutional architectures to modern transformer-based designs. In particular, Vision Transformers (ViTs) have gained traction for their ability to model long-range dependencies using multi-head self-attention and minimal inductive bias [9]. In parallel, the community has revisited large kernel convolutions as a scalable alternative to attention mechanisms, particularly in the context of high-resolution 3D volumetric data [24, 18]. Despite architectural differences, both ViTs and large-kernel CNNs share a central goal: expanding the effective receptive field (ERF) to enable rich spatial context aggregation. However, simply increasing kernel size does not guarantee improved performance. Prior work has shown that naïve enlargement of convolutional filters can result in saturated or degraded accuracy across various segmentation tasks [7, 19]. Unlike ViTs, which adaptively attend to spatial content, standard convolutions rely on static, weight-shared kernels and lack the ability to modulate importance across spatial positions. This limitation prompts our first research question: Can we incorporate spatial priors into large kernel convolutions to improve learning effectiveness?
@@
Recent advances in structural re-parameterization offer a promising direction. Methods such as RepLKNet [7], SLaK [22], and PELK [3] scale kernels to extreme sizes (e.g., $31\times 31$, $51\times 51$, $101\times 101$) by combining parallel branches of “large + small” convolutions into what is referred to as a Constant-Scale Linear Addition (CSLA) block. These parallel paths are merged into a single kernel at inference time, enabling efficient deployment while capturing multi-scale features during training. Interestingly, we observe that CSLA blocks naturally encode spatial learning bias: elements near the kernel center tend to converge faster than those on the periphery. This mirrors diffusion-like gradient propagation in ERFs starting from the center and expanding outward. These observations suggest that convergence dynamics are not uniform across the kernel, but instead spatially structured. This leads to our second question: Can we explicitly model this diffusion pattern as a learnable spatial prior to re-weight kernel element updates during training?
@@
To address this, we first provide a theoretical analysis of the optimization dynamics in CSLA-based re-parameterized convolutions. We show that each branch (e.g., small vs. large kernels) can implicitly operates under a distinct learning rate, leading to element-wise differences in convergence speed. These dynamics correlate with ERF visualizations and share characteristics with spatial frequency patterns in human visual perception [17]. Inspired by this, we propose a novel receptive bias re-parameterization strategy that encodes spatial distance from the kernel center as a spatial bias prior on learning convergence. We implement this as a low-rank modulation mechanism that generates spatial scaling factors for kernel weights, allowing the optimizer to emphasize local versus global regions adaptively for gradient back-propagation.
Building on this insight, we present Rep3D, a 3D convolutional architecture that integrates large kernel convolutions (e.g., $21\times 21\times 21$) with our proposed re-parameterization approach. Unlike prior approaches that rely on multi-branch structures, Rep3D employs a plain and efficient encoder to reduce complexity while preserving representational capacity. We evaluate Rep3D across five challenging volumetric medical segmentation benchmarks and show that it consistently outperforms state-of-the-art transformer- and CNN-based models.Our key contributions are as follows:
@@
Figure 1: (a) Traditional structural re-parameterization methods (e.g., CSLA blocks) re-parameterize small and large kernel convolutions to improve representational capacity, but apply linear optimization with same learning rate across the kernels, demonstrating a faster convergence in local then global. (b) In contrast, Rep3D introduces a learnable spatial bias via a generator network $f_{\theta}$, which modulates each element in the large kernel using a prior based on distance decay. This adaptive modulation enables local-to-global update dynamics aligned with ERF behavior, enhancing both training stability and model performance for 3D volumetric tasks.
@@ 2 Related Work
The Transition from CNN to ViT. Convolutional neural networks (CNNs) have long served as the foundation for medical image segmentation, with the U-Net architecture [27, 5] establishing a dominant encoder–decoder paradigm for dense prediction. Variants such as V-Net [25], UNet++ [33], H-DenseUNet [21], and SegResNet [26] extended this architecture to better suit 2D and 3D modalities, as well as different anatomical contexts. More recently, nnU-Net [14] automated the design of 3D segmentation pipelines with a coarse-to-fine strategy tailored for various medical datasets. However, most CNN-based architectures rely on small kernel sizes, limiting their effective receptive field (ERF) and making it difficult to capture long-range dependencies. Vision Transformers (ViTs), starting with TransUNet [4], brought global attention mechanisms to medical segmentation, allowing models to attend across distant voxels. Models such as UNETR [11], CoTr [29], and LeViT-UNet [30] have shown strong performance, particularly in organ and tumor segmentation. Yet, the quadratic complexity of self-attention poses major bottlenecks for high-resolution volumetric data. To mitigate this, hierarchical transformers such as Swin Transformer [23] introduced localized attention via a sliding window mechanism. Follow-up models like SwinUNETR [10], nnFormer [32], Swin-Unet [2], and SwinBTS [16] adapted this approach to efficiently model multi-scale features. Despite these innovations, transformer-based methods remain computationally expensive and slow to train, particularly for dense 3D segmentation tasks. Meanwhile, depthwise large-kernel CNNs (e.g., ConvNeXt [24]) offer a promising compromise by mimicking ViT-like receptive fields with fewer computational demands. 3D UX-Net [18] and similar designs apply this principle to volumetric medical data, although challenges remain in segmenting fine-scale anatomical structures or multi-scale lesions without additional spatial priors or architectural adaptations.
@@
The integration of Weight Re-parameterization. Structural re-parameterization (SR) has emerged as a powerful paradigm to enhance CNN training without altering inference-time complexity. Models like RepVGG [8] and OREPA [13] employ additional convolution branches (e.g., $1\times 1$ or identity paths) during training to improve gradient flow and feature diversity. These branches are merged into a single convolution kernel post-training, allowing for efficient inference. RepLKNet [7] and SLaK [22] extend this approach to large 2D kernels (e.g., $31\times 31$ and $51\times 51$), increasing the receptive field while maintaining tractable inference cost via kernel decomposition or sparse groups. A complementary line of work focuses on gradient re-parameterization instead of modifying model weights directly. RepOptimizer [6], for example, modifies the back-propagation process by applying learnable scaling to gradient updates, enabling effective training of plain CNNs. These techniques reduce reliance on complex architectural design and have been shown to match or exceed the performance of more intricate networks. While much of the re-parameterization research has focused on 2D natural images, extending these methods to 3D medical imaging presents unique challenges. Volumetric kernels require significantly more parameters, and naive kernel expansion leads to high computational costs and optimization instability. 3D RepUX-Net [19] demonstrate the inital attempt of adapting weight re-parameterization to 3D medical imaging and scale large depthwise kernels with fixed prior context, but still lacks of flexibility on adapting dynamic variation in fine-grained semantics for learning convergence. To bridge this gap, there is growing interest in using spatial priors or effective receptive field modeling to guide re-parameterization for large kernel learning in the 3D setting.
@@ 3 Rep3D
Rep3D rethinks the training dynamics of large-kernel convolution by explicitly embedding spatial bias, derived from effective receptive fields (ERFs), into the optimization process. Motivated by structural reparameterization (SR) and the distinctive gradient behavior observed in ERFs, Rep3D introduces a low-rank, learnable reparameterization that adapts element-wise update behavior across the kernel. We first derive the theoretical equivalence between parallel convolution branches and their single-operator counterparts, showing that a “large + small” convolution block (as in RepLKNet [7]) implicitly assigns spatially varying learning rates. We then translate this insight into a unified formulation and construct a lightweight generator that outputs a convergence-aware modulation mask. The output modulated mask models fine-grained learning dynamics during training, improving both scalability and performance in 3D tasks with large kernel convolution.
@@ 3.1 Variable Learning Convergence in Parallel Branch
As shown in Figure 2, the learning convergence of the large kernel convolution can be improved by either adding up the encoded outputs of parallel branches weighted by diverse scales with SR (RepLKNet [7]) or performing Gradient Reparameterization (GR) by multiplying with constant values (RepOptimizer [6]) in a Single Operator (SO). Inspired by the concepts of SR and GR, we extend the theoretical derivation from RepOptimizer and observe the variable learning rate across branches. We begin by analyzing the CSLA block, a basic two-branch design used in SR-based networks (e.g., RepLKNet [7]). Let $X$ denote the input feature map, and let $W_{L},W_{S}$ be large and small 3D convolution kernels, scaled by fixed positive scalars $\alpha_{L}$ and $\alpha_{S}$, respectively. The output of the CSLA module is:
@@
During training with first-order optimization (i.e. SGD, AdamW) and step size $\lambda$, we apply the stochastic gradient descent rule and update the gradients for the parallel branches as follow:
@@
As the parallel branch architecture updates $W_{L}$ and $W_{S}$ independently:
$W_{L(t+1)}$ $=W_{L(t)}-\lambda_{L}\frac{\partial\mathcal{L}}{\partial W_{L(t)}},\quad W_{S(t+1)}=W_{S(t)}-\lambda_{S}\frac{\partial\mathcal{L}}{\partial W_{S(t)}}.$ (5)
where $\lambda_{L}$ and $\lambda_{S}$ are the learning rate for corresponding branch respectively. Substituting these into the equivalent kernel formulation yields:
$W^{\prime}_{t+1}$ $=\alpha_{L}W_{L(t+1)}+\alpha_{S}W_{S(t+1)}$ (6)
$=\alpha_{L}\left(W_{L(t)}-\lambda_{L}\frac{\partial\mathcal{L}}{\partial W_{L(t)}}\right)+\alpha_{S}\left(W_{S(t)}-\lambda_{S}\frac{\partial\mathcal{L}}{\partial W_{S(t)}}\right)$ (7)
$={\alpha_{L}}W_{L(t)}+{\alpha_{S}}W_{S(t)}-{\lambda_{L}}{\alpha_{L}}{\frac{\partial{\mathcal{L}}}{\partial{W_{L(t)}}}}-{\lambda_{S}}{\alpha_{S}}{\frac{\partial{\mathcal{L}}}{\partial{W_{S(t)}}}}$ (8)
$=W^{\prime}_{t}-{\lambda_{L}}{\alpha_{L}}{\frac{\partial{\mathcal{L}}}{\partial{W_{L(t)}}}}-{\lambda_{S}}{\alpha_{S}}{\frac{\partial{\mathcal{L}}}{\partial{W_{S(t)}}}}$ (9)
From the equation 9, we observe that each branch can be optimized differently with different learning rates toward each kernel and derive with two distinctive scenarios as follow:
$W^{\prime}_{t+1}=\begin{cases}W^{\prime}_{t}-\lambda\left(\alpha_{L}\frac{\partial\mathcal{L}}{\partial W_{L(t)}}+\alpha_{S}\frac{\partial\mathcal{L}}{\partial W_{S(t)}}\right),&\text{if}\quad\lambda_{L}=\lambda_{S}\\ W^{\prime}_{t}-{\lambda_{L}}{\alpha_{L}}{\frac{\partial{\mathcal{L}}}{\partial{W_{L(t)}}}}-{\lambda_{S}}{\alpha_{S}}{\frac{\partial{\mathcal{L}}}{\partial{W_{S(t)}}}},&\text{if}\quad\lambda_{L}\neq\lambda_{S}\end{cases}$ (10)
@@
$\frac{\partial\mathcal{L}}{\partial W_{L(t)}}$ $=\frac{\partial\mathcal{L}}{\partial Y_{\text{CSLA}}}\cdot\frac{\partial Y_{\text{CSLA}}}{\partial W_{L(t)}}=\alpha_{L}\cdot\frac{\partial\mathcal{L}}{\partial Y_{\text{CSLA}}}\cdot\frac{\partial(X\ast W_{L})}{\partial W_{L(t)}},$ (11)
$\frac{\partial\mathcal{L}}{\partial W_{S(t)}}$ $=\frac{\partial\mathcal{L}}{\partial Y_{\text{CSLA}}}\cdot\frac{\partial Y_{\text{CSLA}}}{\partial W_{S(t)}}=\alpha_{S}\cdot\frac{\partial\mathcal{L}}{\partial Y_{\text{CSLA}}}\cdot\frac{\partial(X\ast W_{S})}{\partial W_{S(t)}}.$ (12)
To validate the above theoretical derivation, we perform ablation studies and found out that variable learning rate of each branch (i.e. $\lambda_{S}=0.0006,\lambda_{L}=0.0002$) demonstrates the best performance with stochastic gradient descent. Since $W_{S}$ has a smaller receptive field than $W_{L}$, and $W_{S}$ primarily contributes to the central region of the equivalent kernel $W^{\prime}$, we argue that:
@@
Both the coefficients $\alpha_{L}$ and $\alpha_{S}$ modulate spatially distinct regions (large kernel contributions dominate the periphery, small kernel contributions dominate the central region), the two–branch block demonstrates a learning-rate field of:
@@
where $\lambda_{eff}$ is the effective element-wise learning rate inherited from the two branch-specific updates.

### v2

# Spatially-Adaptive Gradient Re-parameterization for 3D Large Kernel Optimization

@@ Abstract
Large kernel convolutions offer a scalable alternative to vision transformers for high-resolution 3D volumetric analysis, yet naively increasing kernel size often leads to optimization instability. Motivated by the spatial bias inherent in effective receptive fields (ERFs), we theoretically demonstrate that structurally re-parameterized blocks induce spatially varying learning rates that are crucial for convergence. Leveraging this insight, we introduce Rep3D, a framework that employs a lightweight modulation network to generate receptive-biased scaling masks, adaptively re-weighting kernel updates within a plain encoder architecture. This approach unifies spatial inductive bias with optimization-aware learning, avoiding the complexity of multi-branch designs while ensuring robust local-to-global convergence. Extensive evaluations on five 3D segmentation benchmarks demonstrate that Rep3D consistently outperforms state-of-the-art transformer and fixed-prior baselines. The source code is publicly available at https://github.com/leeh43/Rep3D.
Large Kernel Convolution, Weight Re-parameterization, Effective Receptive Field Modeling, Medical Imaging
@@ 1 Introduction
The landscape of medical vision models has evolved rapidly, expanding from early convolutional architectures to modern transformer-based designs. In particular, Vision Transformers (ViTs) have gained traction for their ability to model long-range dependencies using multi-head self-attention and minimal inductive bias (Dosovitskiy et al., 2020). In parallel, the community has revisited large kernel convolutions as a scalable alternative to attention mechanisms, particularly in the context of high-resolution 3D volumetric data (Liu et al., 2022b; Lee et al., 2022). Despite architectural differences, both ViTs and large-kernel CNNs share a central goal: expanding the effective receptive fields (ERFs) to enable rich spatial context aggregation. However, simply increasing kernel size does not guarantee improved performance. Prior work has shown that naive enlargement of convolutional filters can result in saturated or degraded accuracy across various segmentation tasks (Ding et al., 2022b; Lee et al., 2023). Unlike ViTs, which adaptively attend to spatial content, standard convolutions rely on static, weight-shared kernels and lack the ability to modulate importance across spatial positions. This limitation prompts our first research question: Can we incorporate spatial priors into large kernel convolutions to improve learning effectiveness?
@@
Recent advances in structural re-parameterization offer a promising direction. Methods such as RepLKNet (Ding et al., 2022b), SLaK (Liu et al., 2022a), and PELK (Chen et al., 2024) scale kernels to extreme sizes (e.g., $31\times 31$, $51\times 51$, $101\times 101$) by combining parallel branches of “large + small” convolutions into what is referred to as a Constant-Scale Linear Addition (CSLA) block. These parallel paths are merged into a single kernel at inference time, enabling efficient deployment while capturing multi-scale features during training. Interestingly, we observe that CSLA blocks naturally encode spatial learning bias: elements near the kernel center tend to converge faster than those on the periphery. This mirrors diffusion-like gradient propagation in ERFs starting from the center and expanding outward. These observations suggest that convergence dynamics are not uniform across the kernel, but instead spatially structured. This leads to our second question: Can we explicitly model this diffusion pattern as a learnable spatial prior to re-weight kernel element updates during training?
@@
To address this, we first provide a theoretical analysis of the optimization dynamics in CSLA-based re-parameterized convolutions. We show that each branch (e.g., small vs. large kernels) implicitly operates under a distinct learning rate, leading to element-wise differences in convergence speed. These dynamics correlate with ERF visualizations and share characteristics with spatial frequency patterns in human visual perception (Kulikowski et al., 1982). Inspired by this, we propose a novel receptive bias re-parameterization strategy that encodes spatial distance from the kernel center as a spatial bias prior on learning convergence. We implement this as a low-rank modulation mechanism that generates spatial scaling factors for kernel weights, allowing the optimizer to emphasize local versus global regions adaptively for gradient back-propagation. Building on this insight, we present Rep3D, a 3D convolutional architecture that integrates large kernel convolutions (e.g., $21\times 21\times 21$) with our proposed re-parameterization approach. Unlike prior approaches that rely on multi-branch structures, Rep3D employs a plain and efficient encoder to reduce complexity while preserving representational capacity. We evaluate Rep3D across five challenging volumetric medical segmentation benchmarks and show that it consistently outperforms state-of-the-art transformer- and CNN-based models. Our key contributions are as follows:
@@
Figure 1: (a) Traditional structural re-parameterization methods (e.g., CSLA blocks) re-parameterize small and large kernel convolutions to improve representational capacity, but apply linear optimization with the same learning rate across the kernels, demonstrating faster convergence in local regions than global ones. (b) In contrast, Rep3D introduces a learnable spatial bias via a generator network $f_{\theta}$, which modulates each element in the large kernel using a prior based on distance decay. This adaptive modulation enables local-to-global update dynamics aligned with ERF behavior, enhancing both training stability and model performance for 3D volumetric tasks.
@@ 2 Related Work
CNN-based 3D Models: Foundational architectures such as 3D U-Net (Çiçek et al., 2016), V-Net (Milletari et al., 2016), and nnU-Net (Isensee et al., 2021) have played a pivotal role in establishing the standard for volumetric medical image segmentation. These models rely on encoder-decoder designs with dense skip connections, offering a strong balance between spatial resolution and semantic representation. Due to their stability, interpretability, and effectiveness without requiring large-scale pre-training, they remain widely used in clinical and research benchmarks. However, their inherently local receptive fields limit their ability to capture long-range dependencies, motivating subsequent architectural innovations aimed at expanding the effective receptive field (ERF).
Vision Transformer-based and Hybrid Models: To overcome the locality constraints of CNNs, transformer-based architectures such as UNETR (Hatamizadeh et al., 2022b) and SwinUNETR (Hatamizadeh et al., 2022a) introduce global self-attention mechanisms that model distant spatial dependencies more effectively (e.g., follow-up models like nnFormer (Zhou et al., 2021), Swin-Unet (Cao et al., 2021), and SwinBTS (Jiang et al., 2022)). These models encode context across entire volumes through hierarchical token representations, marking a major shift in design philosophy. However, they typically require large-scale pre-training and introduce significant computational complexity due to the quadratic scaling of attention, particularly problematic in 3D volumetric settings. Additionally, their reliance on patch-based tokenization can compromise fine-grained spatial precision—crucial in dense prediction tasks like medical segmentation.
Large Kernel Convolution Networks: A more recent and efficient alternative to transformers involves expanding the ERF through large kernel convolutions, as demonstrated by models such as ConvNeXt (Liu et al., 2022b), 3D UX-Net (Lee et al., 2022), and MedNeXt (Roy et al., 2023). These architectures leverage depth-wise or separable convolutions to approximate global context modeling while preserving the simplicity and inductive biases of convolutional designs. However, studies in 2D vision backbones (e.g., RepLKNet (Ding et al., 2022b) (kernel size: $31\times 31$), SLaK (Liu et al., 2022a) (kernel size: $51\times 51$)) reveal that naively scaling up kernel size leads to saturation or performance degradation in the absence of additional structural guidance. This key insight motivates the design of Rep3D, which augments large 3D kernels with a learnable spatial prior inspired by ERF theory. By explicitly guiding convergence dynamics across kernel elements, Rep3D enables more effective utilization of large kernels, bridging the gap between CNN efficiency and transformer-like contextual modeling.
@@
The Integration of Weight Re-parameterization. Structural re-parameterization (SR) has emerged as a powerful paradigm to enhance CNN training without altering inference-time complexity. Models like RepVGG (Ding et al., 2021) and OREPA (Hu et al., 2022) employ additional convolution branches (e.g., $1\times 1$ or identity paths) during training to improve gradient flow and feature diversity. These branches are merged into a single convolution kernel post-training, allowing for efficient inference. RepLKNet (Ding et al., 2022b) and SLaK (Liu et al., 2022a) extend this approach to large 2D kernels, increasing the receptive field while maintaining tractable inference cost via kernel decomposition or sparse groups. A complementary line of work focuses on gradient re-parameterization instead of modifying model weights directly. RepOptimizer (Ding et al., 2022a), for example, modifies the back-propagation process by applying learnable scaling to gradient updates, enabling effective training of plain CNNs. While much of the re-parameterization research has focused on 2D natural images, extending these methods to 3D medical imaging presents unique challenges. Volumetric kernels require significantly more parameters, and naïve kernel expansion leads to high computational costs and optimization instability. 3D RepUX-Net (Lee et al., 2023) demonstrated the initial attempt of adapting weight re-parameterization to 3D medical imaging and scale large depthwise kernels with fixed prior context, but it still lacks flexibility in adapting dynamic variations in fine-grained semantics for learning convergence. To bridge this gap, there is growing interest in using spatial priors or effective receptive field modeling to guide re-parameterization for large kernel learning in the 3D setting.
@@ 3 Rep3D
Rep3D rethinks the training dynamics of large-kernel convolution by explicitly embedding spatial bias, derived from effective receptive fields (ERFs), into the optimization process. Motivated by structural re-parameterization (SR) and the distinctive gradient behavior observed in ERFs, Rep3D introduces a low-rank, learnable re-parameterization that adapts element-wise update behavior across the kernel. We first derive the theoretical equivalence between parallel convolution branches and their single-operator counterparts, showing that a “large + small” convolution block (as in RepLKNet (Ding et al., 2022b)) implicitly assigns spatially varying learning rates. We then translate this insight into a unified formulation and construct a lightweight generator that outputs a convergence-aware modulation mask, as shown in Figure 2. The output modulated mask models fine-grained learning dynamics during training, improving both scalability and performance in 3D tasks with large kernel convolution.
@@ 3.1 Variable Learning Convergence in Parallel Branch
As shown in Figure 2, the learning convergence of large kernel convolution can be improved by either adding up the encoded outputs of parallel branches weighted by diverse scales with SR (RepLKNet (Ding et al., 2022b)) or performing Gradient Re-parameterization (GR) by multiplying with constant values (RepOptimizer (Ding et al., 2022a)) in a Single Operator (SO). Inspired by the concepts of structural re-parameterization (SR) and gradient re-parameterization, we extend the theoretical derivation from RepOptimizer and observe the variable learning rate across branches. We begin by analyzing the CSLA block, a basic two-branch design used in SR-based networks. Let $X$ denote the input feature map, and let $W_{L},W_{S}$ be large and small 3D convolution kernels, scaled by fixed positive scalars $\alpha_{L}$ and $\alpha_{S}$, respectively. The output of the CSLA module is:
@@
During training with first-order optimization (i.e., SGD, AdamW) and step size $\lambda$, we apply the stochastic gradient descent rule and update the gradients for the parallel branches as follows:
@@
As the parallel branch architecture updates $W_{L}$ and $W_{S}$ independently, the specific updates are:
$W_{L(t+1)}$ $=W_{L(t)}-\lambda_{L}\frac{\partial\mathcal{L}}{\partial W_{L(t)}},$ (5)
$W_{S(t+1)}$ $=W_{S(t)}-\lambda_{S}\frac{\partial\mathcal{L}}{\partial W_{S(t)}},$ (6)
where $\lambda_{L}$ and $\lambda_{S}$ are the learning rates for the corresponding branches. Substituting these into the equivalent kernel formulation yields:
$W^{\prime}_{t+1}$ $=\alpha_{L}W_{L(t+1)}+\alpha_{S}W_{S(t+1)}$
$=\alpha_{L}\left(W_{L(t)}-\lambda_{L}\frac{\partial\mathcal{L}}{\partial W_{L(t)}}\right)$
$\quad+\alpha_{S}\left(W_{S(t)}-\lambda_{S}\frac{\partial\mathcal{L}}{\partial W_{S(t)}}\right)$
$=W^{\prime}_{t}-\lambda_{L}\alpha_{L}\frac{\partial\mathcal{L}}{\partial W_{L(t)}}-\lambda_{S}\alpha_{S}\frac{\partial\mathcal{L}}{\partial W_{S(t)}}.$ (7)
From (7), we observe that each branch can be optimized differently with distinct learning rates toward each kernel, leading to two distinctive scenarios:
$W^{\prime}_{t+1}=\begin{cases}\begin{aligned} &W^{\prime}_{t}-\lambda\Big(\alpha_{L}\frac{\partial\mathcal{L}}{\partial W_{L(t)}}\\ &\quad+\alpha_{S}\frac{\partial\mathcal{L}}{\partial W_{S(t)}}\Big),\end{aligned}&\text{if }\lambda_{L}=\lambda_{S},\\[18.0pt] \begin{aligned} &W^{\prime}_{t}-\lambda_{L}\alpha_{L}\frac{\partial\mathcal{L}}{\partial W_{L(t)}}\\ &\quad-\lambda_{S}\alpha_{S}\frac{\partial\mathcal{L}}{\partial W_{S(t)}},\end{aligned}&\text{if }\lambda_{L}\neq\lambda_{S}.\end{cases}$ (8)
@@
$\frac{\partial\mathcal{L}}{\partial W_{L(t)}}$ $=\frac{\partial\mathcal{L}}{\partial Y_{\text{CSLA}}}\cdot\frac{\partial Y_{\text{CSLA}}}{\partial W_{L(t)}}$
$=\alpha_{L}\cdot\frac{\partial\mathcal{L}}{\partial Y_{\text{CSLA}}}\cdot\frac{\partial(X\ast W_{L})}{\partial W_{L(t)}},$ (9)
$\frac{\partial\mathcal{L}}{\partial W_{S(t)}}$ $=\frac{\partial\mathcal{L}}{\partial Y_{\text{CSLA}}}\cdot\frac{\partial Y_{\text{CSLA}}}{\partial W_{S(t)}}$
$=\alpha_{S}\cdot\frac{\partial\mathcal{L}}{\partial Y_{\text{CSLA}}}\cdot\frac{\partial(X\ast W_{S})}{\partial W_{S(t)}}.$ (10)
To validate the above derivation, we perform ablation studies and find that variable learning rates for each branch (i.e., $\lambda_{S}=0.0006,\lambda_{L}=0.0002$) demonstrate the best performance with stochastic gradient descent. Since $W_{S}$ has a smaller receptive field than $W_{L}$, and $W_{S}$ primarily contributes to the central region of the equivalent kernel $W^{\prime}$, we argue the following:
@@
Because the coefficients $\alpha_{L}$ and $\alpha_{S}$ modulate spatially distinct regions (large kernel contributions dominate the periphery, small kernel contributions dominate the central region), the two-branch block demonstrates an effective learning-rate field of:
@@
where $\lambda_{\mathrm{eff}}$ is the effective element-wise learning rate inherited from the two branch-specific updates.